Measuring growth)  Solarpower Systems earned ​$20 per share at the beginning of the year and paid out ​$8 in dividends to shareholders​ (so, Upper D 0 equals and retained ​$12 to invest in new projects with an expected return on equity of 19 percent. In the​ future, Solarpower expects to retain the same dividend payout​ ratio, expects to earn a return of 19 percent on its equity invested in new​ projects, and will not be changing the number of shares of common stock outstanding.

Calculate the future growth rate for? Solarpower's earnings.
b.If the? investor's required rate of return for? Solarpower's stock is 13 percent?, what would be the price of? Solarpower's common? stock?
c.What would happen to the price of? Solarpower's common stock if it raised its dividends to ?$13 and then continued with that same dividend payout ratio? permanently? Should Solarpower make this? change? ? (Assume that the? investor's required rate of return remains at 13 percent?.)
d.What would happened to the price of? Solarpower's common stock if it lowered its dividends to ?$2 and then continued with that same dividend payout ratio? permanently? Does the constant dividend growth rate model work in this? case? Why or why? not? ? (Assume that the? investor's required rate of return remains at 13 percent and that all future new projects will earn 19 ?percent.
Answer #1

a]

growth rate = retention ratio * ROE of new projects

retention ratio = (retained earnings per share / earnings per share)

growth rate = ($12 / $20) * 19%

growth rate = 11.4%

b]

Price of stock = D1 / (required return - growth rate)

D1 = D0 * (1 + growth rate)

D1 = $8 * (1 + 11.4%) = $8.912

Price of stock = $8.912 / (13% - 11.4%)

Price of stock = $557

c]

growth rate = retention ratio * ROE of new projects

retention ratio = (retained earnings per share / earnings per share)

growth rate = ($7 / $20) * 19%

growth rate = 6.65%

Price of stock = D1 / (required return - growth rate)

Price of stock = $13 / (13% - 6.65%)

Price of stock = $204.72

d]

growth rate = retention ratio * ROE of new projects

retention ratio = (retained earnings per share / earnings per share)

growth rate = ($18 / $20) * 19%

growth rate = 17.1%

Price of stock = D1 / (required return - growth rate)

Price of stock = $2 / (13% - 17.1%)

The constant dividend growth rate model does not work because the growth rate is higher than the required return.

The constant dividend growth rate model does not work when the growth rate is higher than the required return.